OpenStudy (anonymous):

how do you find the rational and real roots? x^7+3x^6-5x^2-3x-5=0

OpenStudy (amistre64):

trial and error.. alot of it lol

OpenStudy (anonymous):

possible rational roots?

OpenStudy (amistre64):

last# ----- helps to narrow it down fisrt#

OpenStudy (anonymous):

Rational root test, fundamental theorem of algebra, synthetic division

OpenStudy (amistre64):

i was thinking of writing a javascript to solve these things ;)

OpenStudy (anonymous):

make a fraction, numerator must divide 5, denom must divide 1, so only choices are \[\pm1, \pm 5\]

OpenStudy (anonymous):

if you can write javascript to find the roots of a poly of degree 7 you would be world famous and galois would have been mistaken.

OpenStudy (amistre64):

no, not all the roots; but the integer and rational roots that could be determined from this lol

OpenStudy (anonymous):

i know i am being silly. you meant to find possible rational roots.

OpenStudy (anonymous):

how do you find the real roots?

OpenStudy (amistre64):

might be able to work it inside out and complete squares to find all roots .... i wonder

OpenStudy (anonymous):

there is no algorithm for finding the roots of a poly of degree seven.

OpenStudy (anonymous):

only 1 - 4. that is it. and the one for 3 is a page and the one for 4 is a wall.

OpenStudy (amistre64):

well, you could do the Newton Rhapsody method to find 1 root and narrow it down ..... perhaps ?

OpenStudy (anonymous):

i would try 1, -1, 5 and -5 and see if any work. then you can use synthetic division to reduce. i can see that 1 does not work since \[f(1)=1+3-5-2-5\]

OpenStudy (amistre64):

make sure you include your zero place holders

OpenStudy (anonymous):

\[f(-1)=-1+3-5+3-5=-1\]

OpenStudy (anonymous):

well we have to find one first.

OpenStudy (anonymous):

if you have a decent calculator check \[f(5)\]

OpenStudy (amistre64):

syntheitc can find them as well; but brute calculating is always a good option lol

OpenStudy (anonymous):

ok i will divide see if i get 0

OpenStudy (watchmath):

Newton-Rhapsody :D :D

OpenStudy (anonymous):

newton rhapsody in blue

OpenStudy (anonymous):

5 doesn't work

OpenStudy (anonymous):

and neither does -5 unless i made a mistake

OpenStudy (amistre64):

-5| 1x^7 +3x^6 +0 +0 +0 -5x^2 -3x -5 0 -5 10 -50 250 -1250 6275 ------------------------------------- 1 -2 10 -50 250 -1255 ...... no mistake lol

OpenStudy (amistre64):

D's sign change tells us what ...... do we include zeros as a sign change?